OpenStudy (anonymous):

x + y + z = -7 2x + y - z =4 -x + 2y + 2z = -17

OpenStudy (amistre64):

x + y + z = -7 2x + y - z =4 -------------- 3x +2y = -3 is a good start

OpenStudy (anonymous):

\[\{x=1,y=-3,z=-5\} \]